# Things we can learn by considering random locally symmetric manifolds

### v1

@@ Abstract
In recent years various results about locally symmetric manifolds were proven using probabilistic approaches. One of the approaches is to consider random manifolds by associating a probability measure to the space of discrete subgroups of the isometry Lie group. The main goals are to prove results about deterministic groups and manifolds by considering appropriate measures. In this overview paper we describe several such results, observing the evolution process of the measures involved. Starting with a result whose proof considered finitely supported measures (more precisely, measures supported on finitely many conjugacy classes) and proceeding with results which were outcome of the successful and popular theory of IRS (invariant random subgroups). In the last couple of years the theory has expanded to SRS (stationary random subgroups) allowing to deal with a lot more problems and establish stronger results. In the last section we shall review a very recent (yet unpublished) result whose proof make use of random subgroups which are not even stationary.
@@ 1 Introduction
The study of discrete subgroups of Lie groups has played a central role in many parts of mathematics since the middle of the previous century, and has been continuously extremely active. In recent years the theory has been boosted by the development of several probabilistic approaches to study these deterministic highly structured objects. In this overview paper I will focus on one of these approaches, namely the consideration of random discrete subgroups. Particularly, I will try to describe my personal journey in the development of the theory.
By a random discrete subgroup we mean a Borel regular probability measure $m$ on the (metrizable, locally compact) space $\text{Sub}_{d}(G)$ of discrete subgroups of $G$. Considering an adjoint semisimple Lie group without compact factors $G$ and letting $X=G/K$ be the associated Riemannian symmetric space, any discrete subgroup $\Lambda\leq G$ corresponds to a locally symmetric orbifold $M=\Lambda\backslash X$ (which is a manifold if $\Lambda$ is torsion free). Thus a random discrete subgroup $m$ corresponds to a random $X$-orbifold.
@@
We start with the result from [GL14]. In that paper A. Levit and the author showed that, counted appropriately, most hyperbolic manifolds are non-arithmetic. More precisely, fixing any dimension $d$ (suppose $d\geq 4$) the number of non-arithmetic complete hyperbolic $d$-manifolds of volume at most $v$ considered up to commensurability is super-exponential as a function of $v$, while the number of arithmetic ones is sub-exponential (see §2 for more precise statements). The proof relies on a model for constructing random hyperbolic manifolds. However, the probability spaces considered in this work are finite, namely we analyze a model for constructing random graphs of manifolds which results on finitely many possibilities corresponding to graphs of bounded size. Therefore, one can argue that this is not a probabilistic proof but a combinatorial one. The corresponding measures $m$ on $\text{Sub}_{d}(G)$ are not finitely supported but rather supported on finitely many conjugacy classes. This is a very primitive form of the general notion of invariant random subgroups (IRS).
An IRS on $G$ is a probability measure on $\text{Sub}(G)$ which is invariant under the $G$-action by conjugation. I believe it is fare to say that the theory of IRS has proven to be much more powerful than the expectation. In particular IRS in semisimple Lie groups turned out to be an extremely significant tool to the study of lattices and their asymptotic invariants. In §3 we shall review some of the basics of IRS in semisimple Lie groups. We shall then demonstrate the power of IRS by presenting two theorems about lattices. The first is the celebrated theorem of Kazhdan and Margulis establishing the lower bound on the co-volume of lattices. In §4 we will present a surprisingly short proof of this classical theorem using a direct analysis of the space $\text{IRS}(G)$ of invariant random subgroups of $G$. The novelty here is not in the result (although the IRS version is not covered by the original proof of Kazhdan and Margulis) but rather in the simplicity and the straightforwardness of the argument.
One of the highlights of the theory of IRS is the Stuck Zimmer rigidity theorem:
@@
In §7 we review the Bengamini–Schramm space $\mathcal{BS}(X)$ associated to a symmetric space $X=G/K$ and its relation with IRS in $G$. The interplay between the geometric point of view naturally associated with $\mathcal{BS}(X)$ and the analytic point of view associated with $\text{IRS}(G)$ is crucial in the proof of the result of the 7 samurai [7’s17]. In §8 we describe the result of [7’s17] concerning the asymptotic geometry of higher rank manifold of finite volume when the volume tends to infinity. The result is that the injectivity radius cannot be bounded when the volume tends to infinity, and in fact that it grows to infinity at ‘almost all points’. More precisely, if $\text{rank}(G)\geq 2$ then $\forall R,\epsilon>0$ there is $V$ such that if $M=\Gamma\backslash X$ has volume $\geq V$ then $\text{vol}(M_{\geq R})/\text{vol}(M)\geq 1-\epsilon$, where $M_{\geq R}$ is the $R$-thick part of $M$ where the injectivity radius is $\geq R$. This geometric result has plenty of applications in the theory of analytic invariants. The most famous application is the convergence of normalized Betti numbers (see [7’s11, ABBG23]), but there are plenty of more refined applications (see [7’s17]). In §9 we will give the proof of the result concerning the Betti numbers in a special case.
While the notion of invariant random subgroups is quite powerful, the same properties that make it so remarkable are also the ones that demonstrate it limitations. For instance, in view of the Stuck–Zimmer rigidity theorem every non-atomic IRS in $G=\text{SL}_{3}({\mathbb{R}})$ is supported on lattices. Therefore IRS are useless in the study of thin subgroups or general discrete subgroups of infinite covolume in $\text{SL}_{3}({\mathbb{R}})$. More generally, as semisimple Lie groups are very far from being amenable, focusing on invariant measures is a huge compromise and restricts by much the scope of problems that can be investigated. For this reason the theory of random subgroups has been expended in the recent years to non-invariant measures.
@@
and let $\nu_{\infty}$ be a weak-* limit. Hopefully by analyzing $\nu_{\infty}$ one can learn about the initial point $\Lambda$. In order to process this philosophy, one should chose the probability measure $\mu$ carefully. In [GLM22] A. Levit, G.A. Margulis and myself have worked out a specific measure $\mu=\mu_{G}$ which posses some wonderful properties. In particular, as shown in [FG23, Theorem 2.2] for every discrete group $\Lambda$, an SRS $\nu_{\infty}$ defined as a weak-* limit of Cesaro averages as above is supported on discrete groups.
@@ 2 Most hyperbolic manifolds are non-arithmetic
In this section revisits my joint work with Arie Levit [GL14].
A celebrated result of Margulis [M84] asserts that finite volume locally symmetric manifolds of rank $>1$ are arithmetic. Corlette [Co92] and Gromov–Schoen [GS92] extended this result to rank one spaces, with the two exceptions of real and complex hyperbolic.
@@
  • Let $Q(x,y,z,w)=x^{2}+y^{2}+z^{2}-\sqrt{2}w^{2}$, let $\mathbb{G}=\mathbb{SO}(Q)$ and let $\Gamma=\mathbb{G}(\mathbb{Z}[\sqrt{2}])$. Then $\Gamma=\mathbb{H}(\mathbb{Z})$ for some algebraic group $\mathbb{H}$ satisfying
@@
In [GPS87] Gromov and Piatetski-Shapiro constructed a non-arithmetic hyperbolic manifold in every dimension. (Prior to that only finitely many examples associated to certain reflection groups were known.) Their idea was to take two arithmetic non-commensurable manifolds $M_{1}$ and $M_{2}$ which can be cut along isometric totally geodesic co-dimension one manifolds, cut them and then glue by identifying the boundaries of the two.
@@
Here is the recipe of the proof of their theorem:
@@
    Let $H\leq G$ be algebraic groups, and $\Gamma\leq G$ an arithmetic subgroup. If $\Gamma\cap H$ is Zariski dense in $H$ then it is arithmetic there. (This is a consequence of the Borel–Harish-Chandra theorem.)
@@
    Let $K$ be a totally real number field. Let $Q_{1},Q_{2}$ be two quadratic forms over $K$ which have signature $(n,1)$ over $\mathbb{R}$ with $n\geq 2$, while all their non-trivial Galois conjugates are positive definite. Let $\Gamma_{i}$ be the arithmetic group in $SO(n,1)$ corresponding to $Q_{i}$ and $M_{i}$ the corresponding hyperbolic orbifold. Then $M_{1}$ and $M_{2}$ are commensurable if and only if $Q_{1}$ is isometric to $\lambda Q_{2}$ for some scalar $\lambda\in K$.
@@
Item $(4)$ implies that in even dimension, the discriminant of the form is invariant under commensurability. This is what stands behind the following:
@@
and let $M_{1},M_{2}$ be the corresponding manifolds (take finite covers to get rid of torsion). Then $M_{1}$ and $M_{2}$ are non-commensurable and admit isometric co-dimension one totally geodesic cuts (along $x_{1}\equiv 0$). Note that $M_{1}$ and $M_{2}$ are non-compact.
@@
These examples thus provide compact and non-compact finite volume non-arithmetic hyperbolic manifolds in any odd dimension. To obtain the even dimensional examples, Gromov and Piatetski-Shapiro took, inside the above example, a co-dimension $1$ sub-manifold which intersect the $N_{i}$’s transversally-orthogonally.
@@
Wang’s proof is non-effective and gives no estimate on $\rho_{n}(v)$. Indeed, Wang’s beautiful argument basically shows that with respect to a certain topology, the set of conjugacy classes of lattices in $SO(n,1)$ is discrete and the co-volume function is proper. I.e. the set of conjugacy classes of lattices of bounded volume is discrete and compact hence finite.
@@
The first motivation for these effective estimates came from theoretical phisics [Ca97, Ca98].
@@
A first lower bound was given by J. Raimbault [^3] Raimbault’s proof is also probabilistic. The reason that he obtained only exponential lower bound while we obtained super-exponential is that he constructed manifolds moduled over circles while we used general regular graphs. Both proofs were inspired by the construction from [7’s12, 7’s20] of IRS in $\text{SO}(n,1)$ which is not contained in any lattice. [R12] who proved $\log C_{n}^{c}(v)\geq\text{Const}\cdot v$. Jointly with A. Levit [CCC] we determined the precise growth:
Theorem 2.12 (G, Levit 2014).
@@
Another number that one may wish to consider is that of torsion free lattices of covolume $\leq v$ considered up to quasi-isometries, denoted $QI_{n}(v)$. However, while all the cocompact lattices in $SO(n,1)$ are quasi-isometric to each other, a beautiful theorem of R. Schwartz [Sch96] says that two non-co-compact lattices in $SO(n,1)$ are quasi isometric iff they are, up to conjugacy, commensurable. I.e.
@@
By a simple counting argument the logarithm of the number of $4$-regular graphs with $v$ vertices is $\approx v\log v$. However, they are all commensurable. By Leighton’s theorem [Lei82], two colored finite graphs are commensurable iff they share the same universal covering colored tree.
@@
Let $X_{1}$ and $X_{2}$ be two submanifolds with totally geodesic boundary and finite volume inside two non-commensurable arithmetic hyperbolic $n$-manifolds. Assume either that $\partial X_{1}$ and $\partial X_{2}$ are compact and $n\geq 3$, or that $\partial X_{1}$ and $\partial X_{2}$ have finite volume and $n\geq 4$.
@@
The proof of this lemma relies on:
@@
The harder case is the even dimension as the discriminant is useless. Instead we address the Legendre and the Hilbert symbols. In the non-compact case we may then use
@@
The proof that the associated manifolds are non-commensurable relies of the following:

### v2

@@ Abstract
In recent years various results about locally symmetric manifolds were proven using probabilistic approaches. One of the approaches is to consider random manifolds by associating a probability measure to the space of discrete subgroups of the isometry Lie group. The main goals are to prove results about deterministic groups and manifolds by considering appropriate measures. In this overview paper, we describe several such results, observing the evolution process of the measures involved. Starting with a result whose proof considered finitely supported measures (more precisely, measures supported on finitely many conjugacy classes) and proceeding with results that were an outcome of the successful and popular theory of IRS (invariant random subgroups). In the last couple of years, the theory has expanded to SRS (stationary random subgroups) allowing to deal with a lot more problems and establish stronger results. In the last section, we shall review a very recent (yet unpublished) result whose proof make use of random subgroups which are not even stationary.
@@ 1 Introduction
The study of discrete subgroups of Lie groups has played a central role in many parts of mathematics since the middle of the previous century, and has been continuously extremely active. In recent years the theory has been boosted by the development of several probabilistic approaches to study these deterministic highly structured objects. In this overview paper, I will focus on one of these approaches, namely the consideration of random discrete subgroups. Particularly, I will try to describe my personal journey in the development of the theory.
Consider an adjoint semisimple Lie group without compact factors $G$. By a random discrete subgroup we mean a Borel regular probability measure $m$ on the (metrizable, locally compact) space $\text{Sub}_{d}(G)$ of discrete subgroups of $G$. Let $K\leq G$ be a maximal compact subgroup and $X=G/K$ the associated Riemannian symmetric space. Any discrete subgroup $\Lambda\leq G$ corresponds to a locally symmetric orbifold $M=\Lambda\backslash X$ (which is a manifold if $\Lambda$ is torsion-free). Thus a random discrete subgroup $m$ corresponds to a random $X$-orbifold.
@@
We start with the result from [GL14]. In that paper A. Levit and the author showed that, counted appropriately, most hyperbolic manifolds are non-arithmetic. More precisely, fixing any dimension $d$ (suppose $d\geq 4$) the number of non-arithmetic complete hyperbolic $d$-manifolds of volume at most $v$ considered up to commensurability is super-exponential as a function of $v$, while the number of arithmetic ones is sub-exponential (see §2 for more precise statements). The proof relies on a model for constructing random hyperbolic manifolds. However, the probability spaces considered in this work are finite, namely, we analyze a model for constructing random graphs of manifolds which results on finitely many possibilities corresponding to graphs of bounded size. Therefore, one can argue that this is not a probabilistic proof but a combinatorial one. The corresponding measures $m$ on $\text{Sub}_{d}(G)$ are not finitely supported but rather supported on finitely many conjugacy classes. This is a very primitive form of the general notion of invariant random subgroups (IRS).
An IRS on $G$ is a probability measure on $\text{Sub}(G)$ which is invariant under the $G$-action by conjugation. I believe it is fair to say that the theory of IRS has proven to be much more powerful than the expectation. In particular, IRS in semisimple Lie groups turned out to be an extremely significant tool to the study of lattices and their asymptotic invariants. In §3 we shall review some of the basics of IRS in semisimple Lie groups. We shall then demonstrate the power of IRS by presenting two theorems about lattices. The first is the celebrated theorem of Kazhdan and Margulis establishing the lower bound on the co-volume of lattices. In §4 we will present a surprisingly short proof of this classical theorem using a direct analysis of the space $\text{IRS}(G)$ of invariant random subgroups of $G$. The novelty here is not in the result (although the IRS version is not covered by the original proof of Kazhdan and Margulis) but rather in the simplicity and the straightforwardness of the argument.
One of the highlights of the theory of IRS is the Stuck–Zimmer rigidity theorem:
@@
In §7 we review the Beniamini–Schramm space $\mathcal{BS}(X)$ associated to a symmetric space $X=G/K$ and its relation with IRS in $G$. The interplay between the geometric point of view naturally associated with $\mathcal{BS}(X)$ and the analytic point of view associated with $\text{IRS}(G)$ is crucial in the proof of the result of the 7 samurai [7’s17]. In §8 we describe the result of [7’s17] concerning the asymptotic geometry of higher rank manifolds of finite volume when the volume tends to infinity. The result is that the injectivity radius cannot be bounded when the volume tends to infinity, and in fact that it grows to infinity at ‘almost all points’. More precisely, if $\text{rank}(G)\geq 2$ then $\forall R,\epsilon>0$ there is $V$ such that if $M=\Gamma\backslash X$ has volume $\geq V$ then $\text{vol}(M_{\geq R})/\text{vol}(M)\geq 1-\epsilon$, where $M_{\geq R}$ is the $R$-thick part of $M$ where the injectivity radius is $\geq R$. This geometric result has plenty of applications in the theory of analytic invariants. The most famous application is the convergence of normalized Betti numbers (see [7’s11, ABBG23]), but there are plenty of more refined applications (see [7’s17]). In §9 we will give the proof of the result concerning the Betti numbers in a special case.
While the notion of invariant random subgroups is quite powerful, the same properties that make it so remarkable are also the ones that demonstrate its limitations. For instance, in view of the Stuck–Zimmer rigidity theorem every non-atomic IRS in $G=\text{SL}_{3}({\mathbb{R}})$ is supported on lattices. Therefore IRS are useless in the study of thin subgroups or general discrete subgroups of infinite covolume in $\text{SL}_{3}({\mathbb{R}})$. More generally, as semisimple Lie groups are very far from being amenable, focusing on invariant measures is a huge compromise and restricts by much the scope of problems that can be investigated. For this reason, the theory of random subgroups has been expanded in the recent years to non-invariant measures.
@@
and let $\nu_{\infty}$ be a weak-* limit. Hopefully by analyzing $\nu_{\infty}$ one can learn about the initial point $\Lambda$. In order to process this philosophy, one should choose the probability measure $\mu$ carefully. In [GLM22] A. Levit, G.A. Margulis, and myself have worked out a specific measure $\mu=\mu_{G}$ which possesses some wonderful properties. In particular, as shown in [FG23, Theorem 2.2] for every discrete group $\Lambda$, an SRS $\nu_{\infty}$ defined as a weak-* limit of Cesaro averages as above is supported on discrete groups.
@@
Acknowledgement: I wish to thank the referee for several good suggestions and for spotting plenty of typos.
@@ 2 Most hyperbolic manifolds are non-arithmetic
A celebrated result of Margulis [M84] asserts that finite volume locally symmetric manifolds of rank $>1$ are arithmetic. Corlette [Co92] and Gromov–Schoen [GS92] extended this result to quaternionic and octonionic hyperbolic spaces. Thus, the only families of finite volume locally symmetric manifolds which may admit non-arithmetic members are the real and complex hyperbolic.
@@
  • Let $Q(x,y,z,w)=x^{2}+y^{2}+z^{2}-\sqrt{2}w^{2}$, let $\mathbb{G}=\mathbb{SO}(Q)$ and let $\Gamma=\mathbb{G}(\mathbb{Z}[\sqrt{2}])$. Then $\Gamma=\mathbb{H}(\mathbb{Z})$ for a $\mathbb{Q}$-group $\mathbb{H}$ satisfying
@@
In [GPS87] Gromov and Piatetski-Shapiro constructed a non-arithmetic hyperbolic manifold in every dimension. (Prior to that only finitely many examples associated with certain reflection groups were known.) Their idea was to take two arithmetic non-commensurable manifolds $M_{1}$ and $M_{2}$ which can be cut along isometric totally geodesic co-dimension one manifolds, cut them, and then glue them by identifying the boundaries of the two.
@@
Here is the recipe for the proof of their theorem:
@@
    Let $H\leq G$ be algebraic groups, and $\Gamma\leq G$ an arithmetic subgroup. If $\Gamma\cap H$ is Zariski dense in $H$, then it is an arithmetic lattice in $H$. (This is a consequence of the Borel–Harish-Chandra theorem. For instance, suppose $\Gamma=G(\mathbb{Z})$ and let $\Delta=\Gamma\cap H$. If $\Delta$ is Zariski dense in $H$, then $H$ is defined over $\mathbb{Q}$ and $\Delta=H(\mathbb{Z})$ is a lattice in $H$.)
@@
    Let $K$ be a totally real number field. Let $Q_{1},Q_{2}$ be two quadratic forms over $K$ which have signature $(n,1)$ over $\mathbb{R}$ with $n\geq 2$, while all their non-trivial Galois conjugates are positive definite. Let $\Gamma_{i}$ be the arithmetic group in $SO(n,1)$ corresponding to $Q_{i}$ and $M_{i}$ the corresponding hyperbolic orbifold. Then $M_{1}$ and $M_{2}$ are commensurable if and only if $Q_{1}$ is isometric to $\lambda Q_{2}$ for some scalar $\lambda\in K$ (see [GPS87, §2.6]).
@@
Item $(4)$ implies that when $n+1$ is even, the discriminant of the form (valued in $k^{*}/(k^{*})^{2}$) is invariant under scalar multiplication and hence well defined on a commensurability class. This is what stands behind the following:
@@
and let $M_{1},M_{2}$ be the corresponding manifolds (take finite covers to get rid of torsion). Then $M_{1}$ and $M_{2}$ are non-commensurable and admit isometric co-dimension one totally geodesic cuts $N_{i}$ (along $x_{2}\equiv 0$). Note that $M_{1}$ and $M_{2}$ are non-compact.
@@
These examples thus provide compact and non-compact finite volume non-arithmetic hyperbolic manifolds in any odd dimension. To obtain the even-dimensional examples, Gromov and Piatetski-Shapiro took, inside the above example, a co-dimension $1$ sub-manifold that intersects the $N_{i}$’s transversally-orthogonally.
@@
Wang’s proof is non-effective and gives no estimate on $\rho_{n}(v)$. Indeed, Wang’s beautiful argument basically shows that with respect to a certain topology, the set of conjugacy classes of lattices in $SO(n,1)$ is discrete and the co-volume function is proper. That is, the set of conjugacy classes of lattices of bounded volume is discrete and compact hence finite.
@@
The first motivation for these effective estimates came from theoretical physics [Ca97, Ca98].
@@
A first lower bound was given by J. Raimbault [^3] Raimbault’s proof is also probabilistic. The reason that he obtained only exponential lower bound while we obtained super-exponential is that he constructed manifolds moduled over circles while we used general regular graphs. Both proofs were inspired by the construction from [7’s12, 7’s20] of IRS in $\text{SO}(n,1)$ which is not contained in any lattice. [R12] who proved $\log C_{n}^{c}(v)\geq\text{Const}\cdot v$. Jointly with A. Levit [GL14] we determined the precise growth:
Theorem 2.12 (Gelander–Levit 2014, [GL14]).
@@
Another number that one may wish to consider is that of torsion-free lattices of covolume $\leq v$ considered up to quasi-isometries, denoted $QI_{n}(v)$. However, while all the cocompact lattices in $SO(n,1)$ are quasi-isometric to each other, a beautiful theorem of R. Schwartz [Sch96] says that two non-co-compact lattices in $SO(n,1)$ are quasi-isometric iff they are, up to conjugacy, commensurable. I.e.
@@
By a simple counting argument, the logarithm of the number of $4$-regular graphs with $v$ vertices is $\approx v\log v$4444444-regular graphs can be thought of as (uncolored) Schrier graphs of $F_{2}$. The latter are in bijection with transitive actions on the set of $v$ elements, i.e. with maps $F_{2}\to\text{Sym}(v)$. Ignoring the transitivity and the coloring issues, there are $v!^{2}$ such maps.
However, they are all commensurable. By Leighton’s theorem [Lei82], two colored finite graphs are commensurable iff they share the same universal covering colored tree.
@@
Let $X_{1}$ and $X_{2}$ be two submanifolds with totally geodesic boundaries and finite volume inside two non-commensurable arithmetic hyperbolic $n$-manifolds. Assume either that $\partial X_{1}$ and $\partial X_{2}$ are compact and $n\geq 3$, or that $\partial X_{1}$ and $\partial X_{2}$ have finite volume and $n\geq 4$.
@@
Making use of the Borel density theorem (lattices are Zariski dense) and the Borel–Harish-Chandra theorem (see Item (1) in the recipe of the proof of GPS above) the proof of the last lemma is reduced to the following geometric fact which is of independent interest:
@@
The harder case is the even dimension as the discriminant is useless. Instead, we address the Legendre and the Hilbert symbols. In the non-compact case, we may then use
@@
The proof that the associated manifolds are non-commensurable relies on the following: